OpenStudy (anonymous):

The circulation ofa newsletter decreased from 3,200 to 2,464. What was the percent of decrease in circulation? A. 129% B.77% C. 2.3% D. 23%

OpenStudy (anonymous):

=(3200-2464)/3200*100 =(736)/32 =23%